OpenStudy (anonymous):

A copper wire of length 2meters, cross sectional area 7.1x10^-6m^2 and Young's modulus 11x10^10N/m^2 has a 200kg load hung on it. What is its increase in length?

OpenStudy (anonymous):

\[\Delta L=\frac {P L}{A E}\] P is the load applied L is the original length A is the cross sectional area E is Young's modulus

OpenStudy (anonymous):

deltaL = 200*2/(7.1x10^-6)*(11x10^10)=.0005?

OpenStudy (anonymous):

when you put in the load for P, remember that force is newtons, not just mass, so do: 200*9.8 not just 200

OpenStudy (anonymous):

thank you that makes sense